a store sells 8 more pencils than pens one day. The cost of the pencil is $0.05 and the cost of the pen is $0.20. If the day's sales of pens and pencils equals $8.90, how many pencils were sold.
I know there is something with x + x+8, but I can't figure out what to do with the different prices. Thanks.

Answer by mananth(16949)

let pens sold be x
pencils sold will be x+8
..
cost of pen = 20 cents
cost of pencils= 5 cents
..
total sales = 890 cents
..
20x+5(x+8)=890
20x+5x+40=890
25x=890-40
25x=850
x=34 pens sold
pencils sold = x+8 = 34+8=42
..
CHECK
34*20+42*5=890